# Echo chamber effects in signed networks

### v1

@@ Results.
For complex contagion, we rely on the SLTM. Agents can now only activate if a sufficiently high number of their neighbors are active themselves. Even with this fundamental difference, our main results regarding the emergence of echo chambers hold. Figure 5 shows that SLTM dynamics again lead to a positive echo chamber effect when $\tau>0$, and an anomalous region in the $\tau<0$ half-plane. Slight differences with the SICM include are the sharper transitions, an increased presence of the anti-echo chamber effect in the $\tau<0$ half-plane, and the persistence of a few anomalous regions under asynchronous activations (Appendix, Sec. D). Overall, these results suggest that the echo chamber phenomenology is robust with regards to the dynamical model employed.
@@
Figure 4: Echo chamber effect with different networks topologies. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.04$ and out-group link density $p_{out}=0.01$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=6$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and $N=250$ agents, except Polblogs with $N=1,222$.
@@
Figure 5: echo chamber effect resulting from a complex contagion process (SLTM) across various topologies. Same networks as in Fig. 4.
Discussion.
@@
Our study makes two key contributions. First, we propose two opinion contagion models, the SICM and the SLTM, that address how information can be framed differently among antagonistic groups. Second, and more importantly, we employ these models to describe the role of structural balance and connectivity in the emergence of echo chambers. Our findings align with previous results connecting polarization with structural balance [31, 22, 34] and connectivity [43]. However, while those studies focused on structural properties, ours is the first to emphasize the impact of these on information spreading. Our results hold for both simple and complex contagion and in a variety of network topologies, as well as in real networks.
This work represents a first step towards incorporating the effect of negative links in the study of echo chambers, with potential extensions to incorporate temporal activity of agents [1], non-Markovian dynamics [62] or homophily [60, 63]. Moreover, the SICM and SLTM show nontrivial regimes and transitions that could be explored more thoroughly. Future works could also consider the effects of structural balance beyond triangles [51, 30], other information transmission biases [60] and examine echo chambers at microscopic and mesoscopic levels [51]. The comparison between theoretical predictions and empirical data (e.g. [6, 57]) is also key to assess the validity of our modelling assumptions. Finally, our results could be applied to mitigate echo chamber effects and polarization in recommendation algorithms [64].
@@ Acknowledgements.
We want to thank the young researchers of the Complex Systems Society for financing a research visit of AV to IFISC through the Bridge Grants program, that served as the primary basis for this work. We also thank Maxi San Miguel for insightful discussions and David Abella for suggesting improvements to the first draft. FDD thanks financial support from MDM-2017-0711-20-2 funded by MCIN/AEI/ 10.13039/501100011033 and FSE, as well as project APASOS (No. PID2021-122256NB-C22). In the early stages of this research, AV was funded by the UK EPSRC grant EP/S022503/1 that supports the Centre for Doctoral Training in Cybersecurity delivered by UCL’s Departments of Computer Science, Security and Crime Science, and Science, Technology, Engineering and Public Policy.
Appendix A Networks generation
To generate signed networks in which to run the dynamical processes, we always follow a three-steps procedure. First, we create an unsigned network through standard generative algorithms, like the Erdös-Rényi or Barabasi-Albert ones. Second, we assign a group to each of the agents in the network. In Erdös-Rényi, Watts-Strogatz and Barabasi-Albert networks, each agent is assigned randomly to one of two groups of the same size, whereas for the Stochastic Block Model and Polblogs, the group labels correspond to the communities existing in the network. Third, we attribute signs to links as follows: links whose endpoints fall within the same group are negative with probability $\eta$ and positive otherwise; whereas links whose endpoints fall in different groups are positive with probability $\eta$ and negative otherwise. The parameter $\eta$ regulates the degree of balance in the network. For $\eta=0$ the network is (perfectly) balanced, for $\eta=1/2$ link signs are uncorrelated with group memberships, while for $\eta=1$ we obtain an antibalanced network. Intermediate values allow us to interpolate between the different degrees of balance. One may then think of $\eta$ as a noise parameter.
To generate networks with a specific triadic degree of balance $\tau$, we make use of the following nonlinear relation, proven in Sec. C.1 below:
${\tau}=(1-2\eta)^{3}$ (3)
Figure 6: Degree of balance as a function of noise, for various networks: ER with density $p=0.1$, ER with density $p=1.0$, BA with average degree $k=3$, and Polblogs [61]. Each point is averaged over 10 random assignments of link signs. Vertical bars indicate one standard deviation from the mean.
Polblogs is a real-life dataset describing connections between political blogs in the US in 2003 [61]. Each node corresponds to a blog and is in one of two communities, Democrat or Republican. The presence of a link between two agents indicates that one mentioned the other. We set group memberships to correspond to community memberships. We restrict ourselves to the largest connected component of the network, which has $1,222$ agents and $19,021$ links.
Table 1 summarizes the main characteristics of the networks. We highlight the fact that Polblogs exhibits both a community structure as the SBM, and local clustering as the WS model. While it is not exactly scale-free like the BA model, its degree distribution is well-fitted by a power-law with exponential cutoff (see [61]), which implies the presence of few hubs and many leaves.
ER BA WS $(q=0.1)$ WS $(q=0.5)$ SBM Polblogs
Average density 0.025 0.024 0.024 0.024 0.025 0.0224
Hubs and leaves - - - -
Communities - - - -
Average modularity - - - - 0.84 0.41
Average clustering coefficient 0.02 0.08 0.45 0.1 0.03 0.32
Table 1: Basic statistics of the networks considered.
Appendix B Simulation details
B.1 Signed Independent Cascade Model (SICM)
The simulation process is as follows. To compute $E$ for Erdös-Rényi networks with density $p$ and degree of balance $\tau$, we first generate $m=20$ such networks via the process described in Sect. A above. Then we perform $2m$ of the SICM with activation rate $\lambda$ on each of these networks, with $m$ simulations seeded in group $a$, and $m$ simulations seeded in group $b$. The number of seeds is chosen to be $10\%$ of the seeding group every time. The remaining agents are initially inactive. At each step, each agent that was activated in the previous step tries to activate independently each of its neighbours. The probability that a given neighbor is activated this way, is the activation rate $\lambda$. If an agent becomes activated through a positive link, it inherits the state ($+1$ or $-1$) of the agent that activated it. If it is activated through a negative link, it instead adopts the opposite state. Once an agent has tried to activate all its inactive neighbors, it ceases to try to activate them in future steps—in other words, each agent tries to activate its neighbors a single time.
For each simulation, we use a breadth-first search algorithm. Contrary to the traditional Independent Cascade, here the specific agent that activates another one matters in the SICM. For example, consider a inactive agent with two neighbors in state $+1$, where one of the neighbors is connected through a positive link and the other through a negative one. The final state of the agent will change depending on which of the agents activates it. Moreover, since the agent cannot update its state once activated, the fist agent that tries to activate has a better chance of propagating its state. To circumvent this issue, we randomize the order in which neighbors try to activate the agent. Finally, we restart the simulation if no agents were activated in a single community. This way, the values of $\rho$ and $E$ are well defined.
We also compare the results with an asynchronous version the process. Whenever an agent becomes activated, each of its neighbors are placed in a queue. Then at each step, an inactive agent is picked at random from the queue and activates with probability $\lambda$. If it succeeds, it adopts the same state as the agent that triggered this activation. This has a clear impact on the results, as the anomalous regimes in the $\tau<0$ region disappear (Figs. 10, 11).
B.2 Signed Linear Threshold Model (SLTM)
We propose the SLTM to adapt the archetypal Threshold Model [45] for signed networks. As for the SICM, all agents are initially in state 00, except a few seeds in state $+1$, who all belong to the same group. At each step, every inactive agent $i$ is selected. The average influence from its neighbours is computed as:
$\phi_{i}(t)=\frac{1}{k_{i}}\sum_{j=1}^{k_{i}}w_{ij}x_{j},$ (4)
where $k_{i}$ is the degree of agent $i$ (independent of the signs of the links), $x_{j}\in\{-1,0,+1\}$ is the state of agent $j$ and $w_{ij}$ the sign of the link between $i$ and $j$. Notice that both $w_{ij}$ and $x_{j}$ can be negative, so $\phi_{i}\in[-1,1].$ We then update the state of $i$ as follows.
$x_{i}\leftarrow\begin{cases}+1&\text{if}\quad\phi_{i}(t)>T,\\ +1&\text{if}\quad\phi_{i}(t)=T\quad\text{with probability }1/2,\\ -1&\text{if}\quad\phi_{i}(t)<-T,\\ -1&\text{if}\quad\phi_{i}(t)=-T\quad\text{with probability }1/2,\\ 0&\text{otherwise}.\end{cases}n$ (5)
The threshold $T\in(0,1)$ quantifies the difficulty with which information flows through the network. Thus, $1-T$ is akin to $\lambda$ in the SICM, as it measures how easily information travels. Note that all activations happen simultaneously each step. In Fig. 12 we compare the results with an asynchronous version the process, where agents are activated one by one in a random order. This has a clear impact on the results, as the anomalous regimes in the $\tau<0$ region (almost) disappear.
The simulation process is similar as for the SICM. To compute $E$, we first generate $m=20$ such networks via the process described in Sect. A above. Then we perform $2m$ of the SLTM with threshold $T$ on each of these networks, with $m$ simulations seeded in group $a$, and $m$ simulations seeded in group $b$. The number of seeds is chosen to be $10\%$ of the seeding group every time.
Appendix C Mathematical proofs
C.1 Triadic balance as a function of noise
We prove that for any network,
$\mathbb{E}[\tau]=(1-2\eta)^{3},$ (6)
where $\tau$ is the degree of triadic balance and $\eta$ is the level of noise as defined in Sec. A above. Given a signed network, consider a triangle $T$ in this network. The triangle has its own degree of balance $\tau(T)$. In the absence of noise ($\eta=0$), we have $\tau(T)=1$. If $\eta>0$, each link in this perfectly balanced version of $T$ is flipped with probability $\eta$.
  • With probability $(1-\eta)^{3}$ no link is flipped, and $T$ stays balanced.
  • With probability $3(1-\eta)^{2}\eta$ exactly one link is flipped, and $T$ becomes antibalanced.
  • With probability $3(1-\eta)\eta^{2}$ two links are flipped, and $T$ stays balanced.
  • With probability: $\eta^{3}$ all three links are flipped, and $T$ becomes antibalanced.
The expected degree of balance of $T$ after the flips is
$\mathbb{E}[\tau(T)]$ $=(1-\eta)^{3}-3(1-\eta)^{2}\eta+3(1-\eta)\eta^{2}-\eta^{3}$
$=(1-2\eta)^{3}.$ (7)
The expected degree of balance $\mathbb{E}[\tau]$ of the whole graph is the average of $\mathbb{E}[\tau(T)]$ over all triangles $T$ after the flips, hence $\mathbb{E}[\tau]=(1-2\eta)^{3}$ as well.
C.2 The echo chamber effect is a second order moment
Let $a,b$ denote two complementary subsets of $V$ representing the two groups in the network, and let $f$ be a vector indicating the group of each agent: $f_{i}=+1$ if $i\in a$, $f_{i}=-1$ if $i\in b$. Similarly, let $I^{\pm}$ represent the set of agents with positive and negative state at the end of the contagion process, and let $x$ be a vector defined as follows: $x_{i}=\pm 1$ if $i\in I^{\pm}$, $x_{i}=0$ otherwise. Furthermore, let $s$ denote the set containing the seeds. Based on these definitions, we can express the fraction of agents in a final state ±$\pm 1$, conditioned on the seeds belonging to group $F_{s}$ and the analyzed agents being active and belonging to group $F_{t}$, as:
$\rho_{F_{s}F_{t}}^{\pm}=\frac{|\left\{i\in F_{t}:x_{i}=\pm 1,s\subset F_{s}\right\}|}{N_{I}(F_{s},F_{t})}$ (8)
where $N_{I}(F_{s},F_{t})$ denotes the final number of active agents belonging to group $F_{t}$. Note that this quantity also depends on the seed group $F_{s}$, as the seed group can affect the efficiency of the contagion process and thus the final number of active agents.
By the law of large numbers, and assuming independent final states (pair approximation [65]), the fraction of active agents in a final state $x_{i}\in\{\pm 1\}$ is approximately equal to the probability of a final state $x_{i}$ in agent $i$ conditioned on $i$ being active and belonging to a given target group $F_{t}$, and the set of seeds belonging to group $F_{s}$:
$\rho_{F_{s}F_{t}}^{\pm}$ =|$=\frac{|\left\{i\in F_{t}:x_{i}=\pm 1,s\subset F_{s}\right\}|}{N_{I}(F_{s},F_{t})}$
$\approx\mathbb{P}[x_{i}=\pm 1|x_{i}\neq 0,s\subset F_{s},i\in F_{t}].$ (9)
To obtain an unconditioned expectation of the number of active agents on a given state, we employ the law of total expectation, $\mathbb{E}_{X,Y}[X]=\mathbb{E}_{Y}[\mathbb{E}_{X}[X|Y]]$, where the subindex indicates the random variable over which the expectation is taken. In this case, the conditional random variable $Y$ corresponds to both the seed and target groups $F_{s}$ and $F_{t}$:
$\mathbb{E}_{x,F_{s},F_{t}}[x_{i}|x_{i}\neq 0]$ $=\mathbb{E}_{F_{s}}[\mathbb{E}_{F_{t}}[\mathbb{E}_{x}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]]]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}\mathbb{P}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}\rho^{x_{i}}_{F_{s}F_{t}}$
$=\frac{1}{4}\left(\rho^{+}_{aa}-\rho^{-}_{aa}+\rho^{+}_{ba}-\rho^{-}_{ba}+\rho^{+}_{ab}-\rho^{-}_{ab}+\rho^{+}_{bb}-\rho^{-}_{bb}\right).$ (10)
where the factor $1/4$ comes from assuming equally likely group membership, so that
$\mathbb{P}(s\subset a)=\mathbb{P}(s\subset b)=\mathbb{P}(i\in a)=\mathbb{P}(i\in b)=1/2.$ (11)
We have thus proved that the expected value of $x_{i}$ (conditioned on being active) is linearly dependent on the variables $\rho_{ab}^{\pm}$, accounting for the sign. We can now do a similar argument to relate the echo chamber effect to these variables. To do so, we first define a variable $c_{si}$ indicating whether the seed and target groups agree:
$c_{si}=\begin{cases}+1\quad\textrm{if}\quad f_{s}=f_{i}\\ -1\quad\textrm{if}\quad f_{s}\neq f_{i},\end{cases}$ (12)
where $i$ is an arbitrary agent, $s$ is the set of seeds and all seeds belong to the same group. Alternatively, we can express $c_{si}$ as a product of the groups of $i$ and $s$: $c_{si}=f_{s}f_{i}$. Note that, since the group assignations can be random, $c_{si}$ is a random variable and we may define a joint probability of $x_{i}$ and $c_{si}$. We now prove that $E$ is a second-order moment of this joint distribution using the methodology employed before:
$\mathbb{E}_{x,F_{s},F_{t}}[x_{i}c_{si}|x_{i}\neq 0]=$
$=\mathbb{E}_{F_{s}}[\mathbb{E}_{F_{t}}[\mathbb{E}_{x}[x_{i}c_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]]]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}f_{i}f_{s}\times$ (13)
×$\quad\quad\quad\times\mathbb{P}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]$
$=\frac{1}{4}\left(\rho^{+}_{aa}-\rho^{-}_{aa}-\rho^{+}_{ba}+\rho^{-}_{ba}-\rho^{+}_{ab}+\rho^{-}_{ab}+\rho^{+}_{bb}-\rho^{-}_{bb}\right)$ (14)
The last expression is the echo chamber effect, so we have proven that $E$ is a second-order moment of the joint distribution of the final states $x_{i}$ and the seed-target group agreeement variable $c_{si}$.
Appendix D Additional figures
Figure 7 shows an example of antibalance in a real-world, signed network from [22]. The agents are members of the U.S. congress and the sign of links connecting them is based on bill co-sponsorships. We see that during the Carter and Reagan administrations (1977-1985), the House of Representatives exhibits a negative degree of balance, and so does the Senate between 1979 and 1983 (96th-97th congress). Thus, in these instances, congressmembers were co-sponsoring bills with colleagues from the other party more than within their own ranks.
Figure 8 gives another perspective on the highly heterogeneous evolution of $E$ with $\lambda$ and $\tau$ in antibalanced networks and under the SICM. Also for antibalanced networks and SICM, Fig. 9 illustrates how the anomalous region changes for simple contagion with a different value of $N$, here $100$ instead of $250$ in the article. The anomalous region has a highly similar shape but is slightly bigger. Finally, Figs. 10 and 11 show that the anomalous region disappear with asynchronous updates in the SICM. They also mostly disappear in the SLTM, except for a tiny one that persists in the Polblogs network (Fig. 12). This last plot was obtained using the SLTM—see Sec. B.2 above for more details.
Figure 7: Degree of Balance in both chambers of the U.S. congress, from 1973 (93rd congress) to 2017 (111st congress).
Figure 8: Echo chamber effect in antibalanced ER networks. (a) $E$ as a function of the $\tau$ for various $\lambda$. (b) $E$ as a function of $\lambda$ for various $\tau$.
Figure 9: Echo chamber effect as a function of network density and activation rate in antibalanced ER networks with $N=100$ agents, for three different $\tau$.
Figure 10: Echo chamber effect in the SICM with asynchronous updates, on ER networks with various densities $p$.
Figure 11: Echo chamber effect in the SICM with asynchronous updates. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.4$ and out-group link density $p_{out}=0.1$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=3$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and a density between $0.022$ and $0.025$, and $N=250$ agents—except Polblogs with $N=1,222$.
Figure 12: Echo chamber effect in the SLTM with asynchronous updates. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.4$ and out-group link density $p_{out}=0.1$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=3$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and a density between $0.022$ and $0.025$, and $N=250$ agents—except Polblogs with $N=1,222$.

### v2

@@ III.2 Robustness
The echo chamber effect emerges when $\tau>0$ for every network, confirming that structural balance favors echo chambers regardless of the topology. Moreover, the anomalous region is also present for every network except the Watts-Strogatz one (Fig. 5f), due to clustering hindering the dominance of even-length paths. For complex contagion, we rely on the SLTM. Agents can now only activate if a sufficiently high number of their neighbors are active themselves. Even with this fundamental difference, our main results regarding the emergence of echo chambers hold. Figure 6 shows that SLTM dynamics again lead to a positive echo chamber effect when $\tau>0$, and an anomalous region in the $\tau<0$ half-plane. Slight differences with the SICM include sharper transitions, an increased presence of the anti-echo chamber effect in the $\tau<0$ half-plane, and the persistence of a few anomalous regions under asynchronous activations (Supplementary Material [57, Sec. III]). Overall, these results suggest that the echo chamber phenomenology is robust with regards to the dynamical model employed.
@@
Figure 5: Echo chamber effect for SICM dynamics, with different networks topologies. (a) ER network with $p=0.025$. (b) Stochastic Block Model (SBM) with in-group link density $p_{in}=0.04$ and out-group link density $p_{out}=0.01$. Community memberships match group memberships. (c) Watts-Strogatz (WS) network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) Barabási-Albert (BA) network with average degree $k=6$. (e) Polblogs network [52]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and $N=250$ agents, except Polblogs with $N=1,222$.
@@
Figure 6: Echo chamber effect $E$ for a signed linear threshold model (SLTM) as a function of the triadic balance $\tau$ and the complementary of threshold parameter, $1-T$, across various topologies. The networks are the same as in Fig. 5.
III.3 Real case study: U.S. congress
Finally, we turn our attention to the study of a collection of real political signed networks between members of the U.S. congress [24]. Each network corresponds to a chamber, the House of Representatives (hereafter “House”) or the Senate, and a congress number between 93rd (1973-1975) and 114th (2015-2017). Positive (negative) edges indicate that legislators co-sponsor significantly more (less) bills than expected under the corresponding null model. The number of congress members varies between 443 and 452 for the House, between 100 and 109 for the Senate. Group memberships correspond to party affiliations, Democrat or Republican. There are a few independent congress members in the data, which we keep in the network but ignore when evaluating the echo chamber effect $E$. We refer the interested reader to the original article of Neal [24] for more details. Importantly, in this network the signs of the edges are obtained from the empirical data, meaning that there is a predefined degree of balance $\tau$ for each network.
To investigate the echo chamber effect in this system, we simulate the SICM on all the networks with varying $\lambda$. The results are presented in Fig. 7, which shows $E$ as a function of $\tau$. We obtain similar results with the SLTM, see the Supplementary Material [57, Sec. III]. The trend is mostly increasing; however, there are three clear instances deviating from this correlation, marked with a blue circle and corresponding to the 93rd, 94th and 95th congresses. This implies that the identification of polarization with echo chambers, while holding true for the majority of cases, can fail under certain structural patterns. To understand the origin of this discrepancy, we have analyzed four characteristics of the network as a function of time: the triadic level of balance $\tau$, the graph density, the ratio of negative to positive edges, and the ratio of Democrats to Republicans (Figure 8). Clearly, the House of Representatives displays several unique structural features during the years where the discrepancy is observed. In particular, we find a very low density of edges and a small ratio of negative edges. We surmise that the combined effect of these two anomalies artificially increase the level of balance and thus the structural polarization, without necessarily impeding the communication between factions. As an aside, note that Fig. 8(a) also exhibits a negative degree of balance during the Carter and Reagan administrations, showing that antibalanced networks arise naturally in real-world systems.
Figure 7: Echo chamber effect $E$ as a function of the degree of balance $\tau$, for SICM dynamics on the U.S. congress networks with various activation probabilities $\lambda$. We perform $M=400$ simulations per data point.
Figure 8: Evolution of the U.S. congress network metrics over time. (a) Degree of balance. (b) Network density. (c) Ratio of negative over positive edges. (d) Ratio of Democrats over Republicans. The grey vertical stripes cover the period from the 93rd to the 96th congresses, during which the echo chamber effect is unexpectedly low with respect to the degree of balance (cf. Fig. 7).
IV Discussion
@@
Our study makes two key contributions. First, we propose two novel contagion models, the SICM and the SLTM, that address how information can be framed differently among antagonistic groups. Second, and more importantly, we employ these models to describe the role of polarization, which we identify with structural balance, in the emergence of echo chambers. Our findings align with previous results connecting polarization with structural balance [33, 24, 36] and connectivity [45]. However, while those studies focused on structural properties, ours is the first to emphasize the impact of these on information spreading. Importantly, we find cases where echo chambers are not correlated with polarization, highlighting that the connection between both phenomena is not always straightforward. Consequently, our findings point out to the dangers of identifying polarization with echo chambers. Our results hold for both simple and complex contagion and in a variety of network topologies, as well as in real-world networks.
This work represents a first step towards incorporating the effect of negative links in the study of echo chambers. We highlighted that some of our results arise as the consequence of the choice of synchronous, discrete time steps for the model. Therefore, the relevance and consequences of different modelling choices should be further investigated. Future works could also consider the effects of structural balance beyond triangles [54, 32], other information transmission biases [64], non-Markovian dynamics [65], or homophily [64, 66]. The theory of multilayer networks could also provide relevant insights into signed spreading phenomena [67]. Moreover, our results could be applied to mitigate echo chamber effects and polarization in recommendation algorithms [68]. Finally, the SICM and SLTM show nontrivial regimes and transitions that could be explored more thoroughly. In this context, an analytic characterization of the contagion processes in signed networks would be highly valuable. Standard frameworks such as Mean Field Theory and the Approximate Master Equation [69] have been adapted to include signed edges in binary state dynamics [70, 41, 71]; however, to the best of our knowledge, no work has considered three-state dynamics with signed edges. We hope that our results motivate researchers to explore this new direction.
@@ Acknowledgements.
We want to thank the young researchers of the Complex Systems Society for financing a research visit of AV to IFISC through the Bridge Grants program, that served as the primary basis for this work. We thank Maxi San Miguel for insightful discussions and David Abella for suggesting improvements to the first draft. We also thank the anonymous reviewers for their precious remarks and suggestions that helped us improve this manuscript. FDD thanks financial support from MDM-2017-0711-20-2 funded by MCIN/AEI/10.13039/501100011033 and FSE, as well as project APASOS (No. PID2021-122256NB-C22). In the early stages of this research, AV was funded by the UK EPSRC grant EP/S022503/1 that supports the Centre for Doctoral Training in Cybersecurity delivered by UCL’s Departments of Computer Science, Security and Crime Science, and Science, Technology, Engineering and Public Policy.